OpenStudy (anonymous):

Can some one tell me if i did it right? I was supposed to solve x^2 - 4x = 12 by completing the square. My work is below.

OpenStudy (anonymous):

x^2 – 4x = 12 (b/2)^2 = (-4/2)^2 = (-2)^2 = 4 x^2 - 4x + 4 = 16 (x - 2)^2 = 16 x - 2 = 4 or x - 2 = -4 x = 6 or x = -2

OpenStudy (anonymous):

thanks(: